OpenStudy (anonymous):

Complete the ordered pairs so that each is a solution of the given equation. x=4+y Ordered pairs; 0, ? -2, ? 4, ?

OpenStudy (anonymous):

Would I take the 0, -2 and 4 and plug it into the x. resulting in 0=4+y, -2=4+y, and 4=4+y?